# Optical and UV Flares from Binary Black Hole Mergers in Active Galactic Nuclei

### v1

@@ II.1 The energy and mass of the disc ejections
which is defined by the bow shock form $R_{\mathrm{b}}(\phi)$ and a maximum angle $\phi_{b}$ limiting the bow shock extension. To define $R_{\mathrm{b}}(\phi)$, we adopt the exact solution of [65] and [66] that describes the bow shock of a spherical wind travelling supersonically through a uniform medium:
$R_{\mathrm{b}}(\phi)=R_{\mathrm{b,0}}\csc\phi\sqrt{3(1-\phi\cot\phi)}.$ (16)
This solution is characterised by the stagnation radius $R_{\mathrm{b,0}}$, which in terms of the quantities of our problem is given by:
$R_{\mathrm{b,0}}=\sqrt{\frac{\dot{M}_{\mathrm{in}}v_{\mathrm{w}}}{4\pi\rho_{\mathrm{d}}v_{\mathrm{k}}^{2}}}.$ (17)
Finally, we set the limiting angle $\phi_{\mathrm{b}}$, such that the back side of the bow shock area is tangent to per-merger cavity, as depicted in Figure 2. Under this consideration, $\phi_{\mathrm{b}}$ is found solving the following equation:
$R_{\mathrm{b}}(\phi_{\mathrm{b}})\cos(\pi-\phi_{\mathrm{b}})=\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}-R_{\mathrm{b,0}}.$ (18)
The way we proceeded to approximate the mass of the plasma ejections given by equations (13)-(18) relies on the formation of a cometary structure due to the interaction of the remnant’s wind and the disc material. We note that such approximation is appropriate when the path of the remnant within the disc is larger than the stagnation radius $R_{\mathrm{b,0}}$. Therefore, in addition to condition (3), we restrict the present analysis to solutions satisfying:
$R_{\mathrm{b,0}}<\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}.$ (19)
From this condition, the limiting angle of integral (15) is restricted within the interval $\phi_{\mathrm{b}}\in(\pi/2,\pi)$.
To obtain an order of magnitude estimation for the upper and lower volumes given by equation 14, we expand the RHS of equation (14) up to first order in $\theta_{\mathrm{k}}$ (which is the limit of interest) and approximate the area $A_{\mathrm{b}}$ by that of a circle of diameter $l_{\mathrm{d}}$ (being $l_{\mathrm{d}}$ the virtual path of the remnant within the intact disc, see Figure 2), giving:
$V_{0}^{+}\approx\frac{\pi}{8}\left(\frac{h_{\mathrm{d}}^{3}}{\theta_{\mathrm{k}}^{2}}-\frac{2h_{\mathrm{d}}^{2}r_{\mathrm{cav}}}{\theta_{\mathrm{k}}}\right),$ (20)
and
$V_{0}^{-}\approx V_{0}^{+}+\frac{2\pi}{8}\frac{h_{\mathrm{d}}^{3}}{\theta_{\mathrm{k}}^{2}},$ (21)
for the upper and lower volumes, respectively. Normalising these volumes by $R_{\mathrm{g}}^{3}$ (where $R_{\mathrm{g}}$ is the gravitational radius of the hosting SMBH), we then estimate:
@@
$\frac{V_{0}^{+}}{R_{\mathrm{g}}^{3}}$ $\approx 1.060\times 10^{6}\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)^{3}\left(\frac{a/R_{\mathrm{g}}}{3000}\right)^{3}\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{-2}$
$-5.028\times 10^{5}\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)^{2}\left(\frac{a/R_{\mathrm{g}}}{3000}\right)^{3}\left(\frac{M_{\mathrm{S}}}{10^{6}\mathrm{M}_{\odot}}\right)^{-1/3}$
×$\times\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{1/3}\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{-1},$ (22)
for the upper expelled volume, and
$\frac{V_{0}^{-}}{R_{\mathrm{g}}^{3}}\approx\frac{V_{0}^{+}}{R^{3}_{\mathrm{g}}}+2.121\times 10^{6}\left(\frac{h_{\mathrm{d}}/a}{0.01}\right)^{3}\left(\frac{a/R_{\mathrm{g}}}{3000}\right)^{3}\left(\frac{\theta_{\mathrm{k}}}{0.1}\right)^{-2},$ (23)
for the lower volume.
@@ II.2 The emission of the disc ejections
Following [67, 62, 24], we approximate the disc ejections as spherical blast waves in homologous expansion [68, 69, 70]. The initial radii of such spheres $R_{0}^{\pm}$ are set so that they contain the volumes of the ejected material before being energised by the remnant (see quation 14):
@@
where the superscript “$\pm$” labels quantities corresponding to the upper (+) and lower (-) ejections. Since the gas density drastically drops outside the disc, we the external radii of the spheres undergoe free expansions:
$R^{\pm}(t^{\prime})=R^{\pm}_{0}+u_{0}^{\pm}(t^{\prime}-t^{\prime}_{0}),$ (25)
where $u_{0}^{\pm}$ is the expansion velocity and $t^{\prime}$ is the elapsed time in the AGN’s frame, being $t^{\prime}=0$ the moment of the coalescence. The radii given by equation 25 is defined for $t^{\prime}\geq t_{0}$, and we estimate the initial expansion time as:
$t^{\prime}_{0}=\Delta t^{\prime}_{\mathrm{cav}}+\Delta t^{\prime}_{\mathrm{HL}}+\Delta t^{\prime}_{\mathrm{h}},$ (26)
where $\Delta t^{\prime}_{\mathrm{cav}}=r_{\mathrm{cav}}/(v_{\mathrm{k}}\cos{\theta_{\mathrm{k}}})$ is the period in which the remnant enters the unperturbed disc, $\Delta t^{\prime}_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{3}$ is the elapsed time before the BH wind is ignited, and $\Delta t^{\prime}_{\mathrm{h}}=\Delta t^{\prime}_{\mathrm{w}}/2$ is the time in which the idealised spheres are formed, which we take as an intermediate point within the wind injection period (see equation 9). To obtain the expansion velocity, we assume equipartition of the total energy of the ejection (see equation 10) among thermal and kinetic energy, i.e., $E^{\pm}_{\mathrm{0,th}}=E^{\pm}_{\mathrm{0,kin}}=E_{0}/2$. Thus, the expansion velocity is taken as:
@@
where $E_{0}$ and $M_{0}^{\pm}$ are given by equations (10) and (13), respectively.
As we show in Appendix A, the gas of the plasma ejections is initially radiation pressure dominated. In this case, the equation of state is approximated as $P_{0}=\frac{1}{3}a_{\mathrm{r}}T_{0}^{4}$, with $a_{\mathrm{r}}$ being the radiation constant and $T_{0}$ the initial temperature. On the other hand, assuming the idealised situation in which the plasma ejections have a homogeneous distribution of matter, their thermal energy and gas pressure can be related as $P_{0}V_{0}=E_{\mathrm{0,th}}(\gamma_{a}-1)$, with $\gamma_{a}=4/3$ being the adiabatic index appropriate for a radiation-dominated gas. Then, the initial temperature of the plasma ejections is estimated as:
@@
To obtain an order of magnitude estimation for typical values of the initial temperatures $T_{0}^{\pm}$, we use the energy and initial volume of the plasma ejection given by equations (12) and (20), keeping the first term (the dominant one) in the latter. Then we obtain:
$T_{0}^{+}\approx\frac{1}{v_{\mathrm{k}}}\left(\frac{\theta_{\mathrm{k}}GM_{\bullet}c}{h_{\mathrm{d}}}\right)^{1/2}\left(\frac{8\eta_{\mathrm{w}}\eta_{\mathrm{in}}\rho_{\mathrm{d}}}{a_{\mathrm{r}}}\right)^{1/4}=1.062\times 10^{5}\mathrm{K}$
×$\times\left(\frac{300\mathrm{kms}^{-1}}{v_{\mathrm{k}}}\right)\left(\frac{0.01}{h_{\mathrm{d}}/a}\right)^{1/2}\left(\frac{10000}{a/R_{\mathrm{g}}}\right)^{1/2}\left(\frac{10^{6}\mathrm{M}_{\odot}}{M_{\mathrm{S}}}\right)^{1/2}$
×$\times\left(\frac{\theta_{k}}{0.1}\right)^{1/2}\left(\frac{M_{\bullet}}{40\mathrm{M}_{\odot}}\right)^{1/2}\left(\frac{\eta_{\mathrm{w}}}{0.005}\right)^{1/4}\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)^{1/4}\left(\frac{n_{\mathrm{i}}}{10^{14}}\right)^{1/4},$ (29)
and $T_{0}^{-}\approx(1/3)^{1/4}T_{0}^{+}$. Since equation (12) underestimates $E_{0}$ and equation (20) over estimates $V_{0}$, the estimated temperature given by (29) is an underestimation for $T_{0}^{\pm}$ when using the full expressions for $E_{0}$ and $V_{0}$ (given by equations 10 and 14, respectively).
For simplicity, we do not calculate the rise of the emission and we consider that photons are trapped until their diffusion timescale meets the expansion dynamical time, which occurs in the period [69, 70]:
@@ II.3 The model parameter space
The EM counterparts discussed in this section are attributed to ejections of plasma driven by a kicked remnant of a BBH merger within an AGN disc. We approximate the thermal emission of such disc eruptions as that of spherical blast waves in homologous expansion [68, 70]. The observed emission corresponds to the photosphere flux given by equation 34, which in the formulation presented here is evaluated by specifying: the distance of the source $D_{\mathrm{L}}$, the mass $M_{\mathrm{S}}$ and accretion rate $\dot{M}_{\mathrm{S}}$ of the AGN’s central engine, the radial distance of the coalescence $a$, the mass of the merger remnant $M_{\bullet}$, the angle $\theta_{\mathrm{k}}$ and velocity $v_{\mathrm{k}}$ of the post-merger gravitational kick, and the inflow and outflow efficiencies $\eta_{\mathrm{in}}$ and $\eta_{\mathrm{w}}$, respectively, of the BH remnant. In what follows, we suggest ranges for the aforementioned parameters, suitable for the assumptions on which the present emission model is built.
The source distance of interest ranges within $D_{\mathrm{L}}\sim[300,3000]$ Mpc, which is compatible with the distances of the BBH coalescences observed during the first three runs of Advanced LIGO [71] and with the distances of AGNs potentially observed with current and forthcoming telescopes. We consider SMBHs with masses within the range of $M_{\mathrm{S}}\sim 5\times 10^{5-7}$ M, as such SMBHs dominate the mass function for AGNs in the local Universe [72, 73]. We are interested in EM counterparts with luminosities comparable to or larger than that of the hosting AGN. Such emissions could occur more effectively in thin discs with relatively low accretion rates, and thus we consider systems with $\dot{m}_{\mathrm{S}}\equiv\dot{M}_{\mathrm{S}}/\dot{M}_{\mathrm{S,Edd}}\sim[0.01,0.1]$. Below $\dot{m}_{\mathrm{S}}\sim 0.01$, the disc experiences a transition to a geometrically thick accretion flow [74], where the present analysis does not apply. The location in the disc where mergers are likely to occur is not well constrained; thus, we consider merger locations ranging within $a\sim[10^{3},5\times 10^{4}]R_{\mathrm{g}}$, consistent with results from previous analyses [35, 75]. The remnant masses of interest fall within $M_{\bullet}\sim[20,80]$ M, which encompasses the bulk of BBH merger remnants seen by previous LVK detections [71]. According to the present analysis, the post-merger kick angle should be small enough to allow sufficient interaction of the remnant with the unperturbed disc region. Thus, we suggest kick angles within $\theta_{\mathrm{k}}\sim[3^{\circ},10^{\circ}]$. The kick velocities in the frame of the coalescence are expected to range within $v_{\mathrm{k}}\sim[100,1000]$ km s-1 [76, 53, 54]. Intuitively, we expect the inflow rate of material onto the kicked BH to be lower than the classical BHL accretion rate (corresponding to the case with no outflow) since BH outflows would prevent some gas inflow. Thus, we suggest inflow efficiencies within $\eta_{\mathrm{in}}\sim[0.05,0.5]$. The outcome of hyper-Eddington inflows onto black holes is poorly studied. Conservatively, here we consider a wind efficiency ranging within $\eta_{\mathrm{w}}\sim[0.001,0.01]$, consistent with previous studies on BH super winds [49, 61]. In addition, a given parameter configuration within the intervals discussed above must fulfil the conditions (3) and (19), i.e.,
$\bm{\max}\left\{R_{\mathrm{b,0}},R_{\mathrm{HL}}\right\}<\frac{h_{\mathrm{d}}}{\sin\theta_{\mathrm{k}}}-\frac{r_{\mathrm{cav}}}{\cos\theta_{\mathrm{k}}}.$ (37)
@@
When employing the present emission model to interpret flares of a particular AGN considering a given GW event, the quantities $D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$, and $M_{\bullet}$ can be constrained by observations. In this case, the possible predicted flares have five free parameters, namely, $a$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$, $\eta_{\mathrm{in}}$, and $\eta_{\mathrm{w}}$.

### v2

@@ II.1 The energy and mass of the disc ejections
with
[…]
where the vertical stratification follows equation 3 and the local speed of sound is constant along the vertical direction. An order of magnitude estimation for the energy $E_{0}$ calculated above can be evaluated considering some mean constant density $\bar{\rho}_{\mathrm{d}}$ and approximating $(s_{\mathrm{M,w}}-s_{\mathrm{m,w}})/\sin\theta_{\mathrm{k}}\approx h_{\mathrm{d}}$, giving:
$E_{\mathrm{0}}$ $\approx 2\pi\eta_{\mathrm{w}}\eta_{\mathrm{in}}\bar{\rho}_{\mathrm{d}}h_{\mathrm{d}}(GM_{\bullet}c)^{2}/v_{\mathrm{k}}^{4}\sim 1.561\times 10^{49}\,\mathrm{erg}$
×$\times\left(\frac{\eta_{\mathrm{w}}}{0.05}\right)\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)\left(\frac{M_{\bullet}}{20\mathrm{M}_{\odot}}\right)^{2}\left(\frac{v_{\mathrm{k}}}{200\,\mathrm{kms}^{-1}}\right)^{-4}$
×$\times\left(\frac{M_{\mathrm{S}}}{10^{7}\,\mathrm{M}_{\odot}}\right)\left(\frac{a/R_{\mathrm{g}}}{10000}\right)\left(\frac{h_{\mathrm{d}}/a}{0.005}\right)\left(\frac{\bar{n}_{\mathrm{d}}}{10^{14}\,\mathrm{cm}^{-3}}\right).$ (15)
The disc ejections are composed of material from the BH wind and the matter it sweeps up from the disc environment. Ultimately, the wind material is taken from the disc itself through the BHL inflow (see Figure 1). Therefore, we consider the mass of the upper and lower ejections to originate in the regions of the disc above and below the BH trajectory (dashed line in Figure 2), up to the point where we assume the wind is quenched (dotted line in Figure 2; see also equation 6). As previously stated, we model the quasi-spherical BH wind as bipolar outflows with a wide opening half-angle $\theta_{\mathrm{w}}$. We then assess the masses of the upper and lower ejections as the material enclosed within the volume generated by the upper and lower triangles of half-opening angles $\theta_{\mathrm{w}}$, translated along the BH trajectory from the disc entry point to the wind quenching point (horizontally bounded by the cavity-disc interface and the vertical dotted line sketched in Figure 2). It can be straightforwardly shown that these upper and lower volumes $V_{0}^{\pm}$ are given by:
[…]
and the corresponding enclosed masses are:
[…]
where
$\begin{bmatrix}A^{+}\\ A_{\triangle}^{+}\\ A_{\triangle}^{-}\\ A^{-}\end{bmatrix}=\frac{\tan\theta_{\mathrm{w}}}{\tan\theta_{\mathrm{k}}}\begin{bmatrix}(s_{\mathrm{M,w}}-s_{\mathrm{m}})(2s-s_{\mathrm{M,w}}-s_{\mathrm{m}})\\ (s-s_{\mathrm{m}})^{2}\\ (s_{\mathrm{M,w}}-s)^{2}\\ (s_{\mathrm{M,w}}-s_{\mathrm{m}})(s_{\mathrm{m}}+s_{\mathrm{M,w}}-2s)\end{bmatrix}.$ (20)
We limit the upper and lower volumes vertically at ±$\pm s_{\mathrm{M}}=\pm 2h_{\mathrm{d}}$. This choice for the upper and lower limits is motivated by the fact that at such heights, the disc density has dropped substantially (to $\sim 10\%$ of the mid-plane density), and at the same time such vertical heights are attained by the wind shock front in shorter time scales compared to the total crossing time of the remnant (this is further detailed by equation II.2 and Figure 10). Summarising, the characterictic heights employing to estimate mass and energy stored in the disc eruption (see equations 13, 16-19) are $s_{\mathrm{m}}=\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}$, $s_{\mathrm{m,w}}=\tan\theta_{w}r_{\mathrm{cav}}+\sin\theta_{\mathrm{k}}R_{\mathrm{HL}}$, $s_{\mathrm{M,w}}=h_{\mathrm{d}}$, $s_{\mathrm{M}}=2h_{\mathrm{d}}$.
An order of magnitude estimation for the volumes $V_{0}^{\pm}$ can be obtained approximating them as half of the cylinder’s volume of diameter $\frac{h_{\mathrm{d}}}{\tan\theta_{\mathrm{k}}}-r_{\mathrm{cav}}$ and height $s_{\mathrm{M,w}}=2h_{\mathrm{d}}$ (see Figure 2):
$V^{\pm}_{0}\approx\frac{\pi}{2}\left(\frac{h_{\mathrm{d}}}{\tan\theta_{\mathrm{k}}}-r_{\mathrm{cav}}\right)^{2}h_{\mathrm{d}}\approx\frac{\pi h_{\mathrm{d}}^{3}}{2\theta_{\mathrm{k}}^{2}},$ (21)
where the second approximation holds for small angles $\theta_{\mathrm{k}}$, which is the case of interest. The last expression in equation 21, clearly over-estimates the volumes given by equations 16-17, but can be useful for a quick order or magnitude evaluation. Normalising the simplified estimation 21 by $R_{g}^{3}$ ($R_{\mathrm{g}}=GM_{\mathrm{S}}/c^{2}$, the gravitational radius of the central SMBH) one obtains:
@@
$\frac{V_{0}}{R_{\mathrm{g}}^{3}}\approx 6.446\times 10^{6}\left(\frac{10^{\circ}}{\theta_{\mathrm{k}}}\right)^{2}\left(\frac{h_{\mathrm{d}}/a}{0.005}\right)^{3}\left(\frac{a/R_{\mathrm{g}}}{10^{4}}\right)^{3}.$ (22)
@@ II.2 The emission of the disc ejections
Following [74, 71, 25], we approximate the disc ejections as spherical blast waves in homologous expansion [75, 76, 77]. The initial radii of such spheres $R_{0}^{\pm}$ are set so that they contain the volumes of the ejected material (see quations 16-17):
@@
where the superscript “$\pm$” labels quantities corresponding to the upper (+) and lower (-) ejections. Since the gas density drastically drops outside the disc, we assume the external radii of the spheres to undergo free expansions:
$R^{\pm}(t^{\prime})=R^{\pm}_{0}+u_{0}^{\pm}(t^{\prime}-t^{\prime\pm}_{0}),$ (24)
where $u_{0}^{\pm}$ is the expansion velocity and $t^{\prime}$ is the elapsed time in the AGN’s frame, being $t^{\prime}=0$ the moment of the coalescence. The radii given by equation 24 is defined for $t^{\prime}\geq t^{\prime\pm}_{0}$, and we estimate the initial expansion time as:
$t^{\prime\pm}_{0}=\Delta t^{\prime}_{\mathrm{cav}}+\Delta t^{\prime}_{\mathrm{HL}}+\Delta t^{\prime\pm}_{\mathrm{bo}},$ (25)
where $\Delta t^{\prime}_{\mathrm{cav}}=r_{\mathrm{cav}}/(v_{\mathrm{k}}\cos{\theta_{\mathrm{k}}})$ is the period in which the remnant enters the unperturbed disc, $\Delta t^{\prime}_{\mathrm{HL}}=2GM_{\bullet}/v_{\mathrm{k}}^{3}$ is the elapsed time before the BH wind is ignited, and $\Delta t^{\prime\pm}_{\mathrm{bo}}$ is the time in which the disc material is expelled above en bellow the disc. We estimate the $\Delta t^{\prime\pm}_{\mathrm{bo}}$ period as the time the BH wind takes to reach the upper and lower boundaries ±$\pm s_{\mathrm{M}}$. Employing the solution of Waever et al. [78] for the evolution of a stellar wind bubble $R_{w}=(\frac{250}{308\pi})^{1/5}L_{\mathrm{w}}^{1/5}\rho_{\mathrm{ext}}^{-1/5}t^{3/5}$ and taking the density of the external medium $\rho_{\mathrm{ext}}$ as the maximum density within the upper and lower break out distances, this results in:
$\Delta t^{\prime\pm}_{\mathrm{bo}}=$ $\,v_{\mathrm{k}}\left[\frac{250}{77}\eta_{\mathrm{w}}\eta_{\mathrm{in}}(GM_{\bullet}c)^{2}\right]^{-1/3}$
×$\times\begin{cases}(s_{\mathrm{M}}-s_{\mathrm{m,w}})^{5/3},\\ \\ (\rho_{\mathrm{c}}/\rho_{\mathrm{m,w}})^{1/3}(s_{\mathrm{M}}+s_{\mathrm{m,w}})^{5/3}.\end{cases}$ (26)
When the BH wind is quenched, two ejections of disc material expand above and bellow the disc. To obtain the expansion velocity of such expelled material, we assume equipartition of their total energy (see equation 13) among thermal and kinetic energy, i.e., $E^{\pm}_{\mathrm{0,th}}=E^{\pm}_{\mathrm{0,kin}}=E_{0}/2$. Thus, the expansion velocity of the ejected material is taken as:
@@
where $E_{0}$ and $M_{0}^{\pm}$ are given by equations (13) and (18-19), respectively.
As we show in Appendix B, the gas of the plasma ejections is initially radiation pressure dominated. In this case, the equation of state is approximated as $P_{0}=\frac{1}{3}a_{\mathrm{r}}T_{0}^{4}$, with $a_{\mathrm{r}}$ being the radiation constant and $T_{0}$ the initial temperature. On the other hand, assuming the idealised situation in which the plasma ejections have a homogeneous distribution of matter, their thermal energy and gas pressure can be related as $P_{0}V_{0}=E_{\mathrm{0,th}}(\gamma_{a}-1)$, with $\gamma_{a}=4/3$ being the adiabatic index appropriate for a radiation-dominated gas. Thus, we calculate the initial temperature of the plasma ejections as:
@@
To illustrate an order of magnitude estimation for typical values of the temperature $T_{0}^{\pm}$, we the use energy and initial volume of the plasma given by equations 15 and 21, giving:
$T_{0}^{+}\approx\frac{1}{v_{\mathrm{k}}}\left(\frac{\theta_{\mathrm{k}}GM_{\bullet}c}{h_{\mathrm{d}}}\right)^{1/2}\left(\frac{2\eta_{\mathrm{w}}\eta_{\mathrm{in}}\rho_{\mathrm{d}}}{a_{\mathrm{r}}}\right)^{1/4}=1.184\times 10^{5}\,\mathrm{K}$
×$\times\left(\frac{200\,\mathrm{kms}^{-1}}{v_{\mathrm{k}}}\right)\left(\frac{0.005}{h_{\mathrm{d}}/a}\right)^{1/2}\left(\frac{5000}{a/R_{\mathrm{g}}}\right)^{1/2}\left(\frac{10^{7}\,\mathrm{M}_{\odot}}{M_{\mathrm{S}}}\right)^{1/2}$
×$\times\left(\frac{\theta_{k}}{10^{\circ}}\right)^{1/2}\left(\frac{M_{\bullet}}{20\,\mathrm{M}_{\odot}}\right)^{1/2}\left(\frac{\eta_{\mathrm{w}}}{0.05}\right)^{1/4}\left(\frac{\eta_{\mathrm{in}}}{0.1}\right)^{1/4}\left(\frac{n_{\mathrm{i}}}{10^{14}\,\mathrm{cm}^{-3}}\right)^{1/4}.$ (29)
For simplicity, we do not calculate the rising emission produced by the disc eruptions and assume that photons remain trapped until their diffusion timescale matches the expansion dynamical time, which occurs over the period [76, 77]:
@@ II.3 The model parameter space
The EM counterparts discussed in this paper are attributed to ejections of plasma driven by a kicked remnant of a BBH merger within an AGN disc. We approximate the thermal emission of such disc eruptions as that of spherical blast waves in homologous expansion [75, 77]. The observed emission corresponds to the photosphere flux given by equation 34, which in the formulation presented here is evaluated by specifying: the distance of the source $D_{\mathrm{L}}$, the mass $M_{\mathrm{S}}$ and accretion rate $\dot{M}_{\mathrm{S}}$ of the AGN’s central engine, the radial distance of the coalescence $a$, the mass of the merger remnant $M_{\bullet}$, the angle $\theta_{\mathrm{k}}$ and velocity $v_{\mathrm{k}}$ of the post-merger gravitational kick, the BH wind opening angle $\theta_{\mathrm{w}}$, and the inflow and outflow efficiencies $\eta_{\mathrm{in}}$ and $\eta_{\mathrm{w}}$, respectively, of the BH remnant. In what follows, we suggest ranges for the aforementioned parameters, suitable for the assumptions on which the present emission model is built.
The source distance of interest ranges within $D_{\mathrm{L}}\sim[300,3000]$ Mpc, which is compatible with the distances of the BBH coalescences observed during the first three runs of Advanced LIGO [79] and with the distances of AGNs potentially observed with current and forthcoming telescopes.
We consider SMBHs with masses within the range of $10^{6-8}$ M as such SMBHs dominate the mass function for AGNs in the local Universe [80, 81]. We are interested in EM counterparts with luminosities comparable to or larger than that of the hosting AGN. Such emissions could occur more effectively in thin discs with relatively low accretion rates, and thus we consider systems with $\dot{m}_{\mathrm{S}}\equiv\dot{M}_{\mathrm{S}}/\dot{M}_{\mathrm{S,Edd}}\sim[0.005,0.1]$. Below the lower limit of this range the disc experiences a transition to a geometrically thick accretion flow [82], limiting the application of the present analysis.
The location in the disc where mergers are likely to occur is not well constrained; thus, we consider merger locations ranging within $a\sim 5\times 10^{2-4}R_{\mathrm{g}}$, consistent with results from previous analyses [36, 83]. The remnant masses of interest fall within $M_{\bullet}\sim[10,150]$ M, which encompasses the inferred remnant masses of previous LVK detections [79].
According to the present analysis, the post-merger kick angle should be small enough to allow sufficient interaction of the remnant with the unperturbed disc region. Thus, we suggest kick angles within $\theta_{\mathrm{k}}\sim[5^{\circ},15^{\circ}]$. The kick velocities in the frame of the coalescence could be as large as $v_{\mathrm{k}}\sim 2000$ km s-1 [84, 57, 59, 60]. Nevertheless, kicks within the previous $\theta_{\mathrm{k}}$ range could be limited to $v_{\mathrm{k}}\sim[100,400]$ km s-1 (assuming a merger with orbital plane parallel to the disc plane, see Section IV), which is the range we consider here.
Recent hydrodynamical simulations of BBHs embedded within AGN discs conducted by [39] and [40], exhibit inflows onto the binaries of about 8% of the classical Bondi rate. However, the nature of the remnant’s inflow analysed here differs from these previous studies as we examine accretion onto a single BH moving at supersonic speeds (then undergoing “wind accretion”) and producing outflows. Thus, for the inflow efficiency parameter, we propose the range of $\eta_{\mathrm{in}}\sim[0.05,0.5]$, which encompasses the results of [39] and [40].
The outcome of hyper-Eddington inflows onto black holes is poorly studied. Conservatively, here we consider a wind efficiency ranging within $\eta_{\mathrm{w}}\sim[0.01,0.1]$, consistent with previous studies on BH super winds [53, 67]. In addition, a given parameter configuration within the intervals discussed above must fulfil the conditions (4) and (7), i.e.,
$\bm{\max}\left\{R_{\mathrm{HL}}\,,\,\tan\theta_{\mathrm{k}}r_{\mathrm{cav}}+\sin\theta_{\mathrm{k}}R_{\mathrm{HL}}\right\}<h_{\mathrm{d}}.$ (37)
@@
When employing the present emission model to interpret flares of a particular AGN considering a given GW event, the quantities $D_{\mathrm{L}}$, $M_{\mathrm{S}}$, $\dot{M}_{\mathrm{S}}$, and $M_{\bullet}$ can be constrained by observations. In this case, the possible predicted flares have six free parameters, namely, $a$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$, $\theta_{\mathrm{w}}$, $\eta_{\mathrm{in}}$, and $\eta_{\mathrm{w}}$.